Fermi National Laboratory
 

Accelerator Update for March 23, 2009 - March 25, 2009

 

Monday March 23

The day shift began with Operations monitoring the Tevatron (TeV) store 6917 (122.95E30) and the Recycler stash (237.6E10), with the Antiproton Source stacking (13.0E10), with Linac experts tuning to bring intensities up, with the Booster and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

Linac intensities returned to normal at 8:47 AM.

At 9 AM, a Linac quadrupole power supply (QPS113) tripped off. It reset.

Operators began preparing the Switchyard enclosures C, D, E, and J for access at 9:25 AM.

At 9:26 AM, a Booster RF station (BRF4) tripped off. It reset.

At 9:29 AM, BRF9 tripped off. It reset.

At 9:38 AM, BRF9 and 15 tripped off. They reset.

At 9:57 AM, Operations reported that sparking had caused BRF4 to trip off several more times.

At 10:17 AM, Operations reported that all of Booster was down.

At 10:24 AM, a Booster expert reported that there was no 480-volt supply for the East or West Booster galleries. The Duty Electrician was called.

The Duty Electrician reported at 10:56 AM that he had brought the Booster West Gallery 480-volts on. Booster technicians were bringing the RF stations on.

The Duty Electrician reported at 11:11 AM that the brought the Booster East Gallery 480-volts on. Booster technicians brought the RF stations back on.

Booster experts reported finding a bug in the software used for tuning the bias curve at 11:29 AM. They began working on a fix.

Booster resumed sending beam at 11:52 AM, but with a lot of sparking RF stations.

Operations terminated store 6917 at 11:53 AM.

At 11:57 AM, Operations reported that the Booster RF was sparking so much that the RF system went down. RF experts investigated and found a problem inside the West Anode Power Supply (WAPS). They began repairs.

Operations prepared CDF for access at 12:08 PM.

Operations prepared D0 for access at 1 PM.

Booster experts reported at 1:06 PM that they had completed work on WAPS. Operations resumed sending beam from Booster at lower intensity while the RF expert watched.

At 1:27 PM, BRF13 tripped off. It reset.

Operations reported at 1:52 PM that they were preparing the Meson enclosures M01 – 5 for access.

At 1:55 PM, BRF7 tripped off. It reset.

Operations reported at 1:58 PM that the Booster was still sparking, but that experts were still tuning.

At 2:04 PM, Operations reported that Linac was not delivering any beam. Experts investigated.

D0 came out of access at 2:07 PM.

Operations reported at 2:28 PM that the Linac beam had returned. It took a while because of a Linac controls problem that was stopping them from seeing that the I- Source was tripping off.

CDF came out of access at 2:39 PM.

Operations put the TeV through a dry squeeze at 3:04 PM.

Operations reported at 3:33 PM that the Recycler had stashed antiprotons four times during the day shift.

Operations put the TeV through a wet squeeze at 3:49 PM.

Operations reported at 4 PM that the TeV orbits looked good.

Operations put the TeV into shot setup at 4:12 PM.

Operations established store 6920 at 5:30 PM with an initial luminosity of 284.04E30.

At 5:48 PM, BRF6 tripped off for the first of five times in about two hours. It reset.

At 8:25 PM, Operations reported that store 6920 aborted. There was no quench. A TeV power supply (T:QFA4) was the only device on the alarm screen. Operations called in experts.

Power supply experts reported at 10:12 PM that T:QFA4 was back online. They replaced the power drawer.

Operations put the TeV through a dry squeeze at 10:22 PM. T:QFA4 followed its reference through the squeeze.

A Booster expert reported at 10:26 PM that the Klystron Debuncher was having temperature phase problems. Operations called in an expert.

Operations reported at 11:53 PM that the Recycler had stashed six times during the evening shift.

Tuesday March 24

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (55.6E10), with Pbar stacking (35.5E10), and with NuMI and MiniBooNE taking beam.

At 12:25 AM, LRF3 tripped off for the first of two times in two minutes. It reset.

Operations established store 6922 at 12:36 AM with an initial luminosity of 170.02E30.

At 12:42 AM, T:QFA4 tripped off and store 6922 aborted. There was no quench. Operations called in power supply experts.

At 1:11 AM, LRF3 tripped off for the first of two times in 18 minutes. It reset.

At 1:47 AM, NuMI tripped off on a MI RF indication, but no RF stations had alarmed. It reset.

At 4:51 AM, BRF6 tripped off. It reset.

Operations put the TeV through a dry squeeze at 5:23 AM.

At 5:24 AM, LRF2 tripped off. It reset.

TeV power supply experts reported at 5:43 AM that T:QFA4 had tripped on a circulating current overcurrent. They replaced the auxiliary power and regulation units and lowered the circulating current.

At 6:36 AM, Operations reported that the Booster Gradient Magnet Power Supply had tripped off. It reset.

Operations put the TeV through a wet squeeze at 7:05 AM.

The Neutron Therapy Facility (NTF) reported at 7:38 AM that they were beginning treatment on their first of four patients for the day.

Operations put the TeV through a wet squeeze at 8:02 AM.

Operations reported at 8:08 AM that the Recycler had stashed seven times during the midnight shift.

Operations put the TeV into shot setup at 8:26 AM.

Operations established store 6926 at 9:35 AM with an initial luminosity of 198.43E30.

MI experts began a MI-8 study at 10:12 AM. MI experts were also looking at beam positions.

At 11:03 AM, Operations reported that the last set of antiproton transfers were bad. Operators investigated and discovered that the 2.5 MHz cavity bias power supply had tripped off on a water turbine indication. Experts reported that an undocumented cable had been disconnected. This problem was fixed so that it would not happen again.

At 11:47 AM, MI was suffering from bad beam positions. Experts reported that this was due to the Booster beam. Booster experts investigated.

NTF reported at 1:49 PM that they had finished treating their last patient for the day.

At 1:53 PM, Operations reported that the Linac beam would be down for about 10 minutes for an I- Source access to reset a refrigerator.

Operations reported at 4:02 PM that the Recycler had stashed four times during the day shift.

A MI expert began a study at 4:56 PM.

Operations reported at 7:13 PM that they had tried to raise the intensity to MiniBooNE, but it caused too many beam position problems. A Proton Source expert investigated.

At 8:51 PM, a TeV sector E4 air compressor tripped off. Operators reset it.

At 9:52 PM, BRF13 tripped off for the first of two times in seven minutes. It reset.

At 9:59 PM, BRF6 and 16 tripped off. They reset.

At 11:14 PM, a Booster vacuum pump tripped off for the first of two times in one minute. It reset.

At 11:46 PM, BRF11 tripped off for the first of two times in two minutes. It reset.

Operations reported at midnight that the Recycler had stashed eight times during the evening shift.

Wednesday March 25

The midnight shift began with Operations monitoring store 6926 (49.35E30) and the stash (308.7E10), with Pbar stacking (36.7E10), and with NuMI and MiniBooNE taking beam.

Operations terminated store 6926 at 1:46 AM.

Operations put the TeV into shot setup at 2:04 AM.

Operations established store 6927 at 3:31 AM with an initial luminosity of 275.78E30.

At 3:34 AM, the Pbar sweeper power supply tripped off. It reset.

At 6:50 AM, the P1 transfer line suffered a vacuum cascade. Operators closed the surrounding vacuum valves and the vacuum began recovering.

NTF reported at 7:27 AM that they were beginning to treat their first of three patients for the day.

At 7:46 AM, LRF4 tripped off. It reset.

Operations reported at 8:04 AM that the Recycler had stashed six times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/25/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory